# Gauging or extending bulk and boundary conformal field theories: Application to bulk and domain wall problem in topological matter and their descriptions by (mock) modular covariant

### v2

@@ VI Acknowledgement
We thank Rui-Dong Zhu for introducing the work [103] with us about six years ago and Yuji Tachikawa and Yunqin Zheng for a fruitful collaboration closely related to new BCFTs in this manuscript. We also thank Urei Miura, Shumpei Iino, Kohki Kawabata, Tokiro Numasawa, Steven Simon, Masahiko G. Yamada, and Yuan Yao for the interesting discussions. We thank Bo Yang and Guangyue Ji for the related collaborations and for notifying us of the recent progress on the studies of the HLL in FQHEs and for teaching us the basic notions in FQHE through their lectures and tutorials. We also thank Hao Gang for notifying the work[118] for us. We acknowledge the support from NCTS and CTC-NTHU.
@@ Appendix A Insertion of topological defect and noninvertible chiral gauging
However, it is known that the $\mathcal{D}_{\psi}\sigma\overline{\sigma}$ corresponds to the disorder operator. This has been established in the study of fermionic models and this counting is known as fermionic zero modes. Hence the interpretation of objects after the contraction needs further careful discussion, whereas they provide the correct partition functions.
In the above discussions, one can see the correspondence between $\phi$ and $\mathcal{D}$ evidently. We stress that the formal chiral and antichiral decomposition of the bulk fields $\Phi$ is not a tensor product. This is called the Deligne product [174] verified by “abstract nonsense" terminologically, but still, the formal decomposition is correct up to the chiral and antichiral decomposition of the partition function. Hence one can trust the discussion on the amplitude in Section. II. As a summary, we note the relevant set of objects and identifications,
@@
The set $\{\mathcal{D}_{\alpha}\}$ corresponds to the chiral twist following modern literature on Fradkin-Kadanoff transformations[335]. Hence one can implement the antichiral twist under the following identification,
@@
In successive discussions, we introduce the modular $S$ transformation to study the symmetry $\{\mathcal{Q}\}$. Here, we assume that the modular $S$ transformation does not change the total Hilbert space. Hence the modular $S$ tarnsformation is a mapping $S:\{\phi\}\times\{\overline{\phi}\}\rightarrow\{\Phi\}$. This can be achieved by constructing the modular $S$ matrix for the chiral and antichiral fields and applying the projection operator to the original Hilbert space. In other words, the construction of the original theory is fundamentally important. This choice of projection may become nontrivial when assuming the existence of extended chiral objects. For example, if one studies an orbifolded model, such as the $D$ series model, the object in the $A$ series model appears as such. Moreover, in the recent studies on Majorana and the parafermionic model, this extension of the Verlinde line also appears. Hence the representation $S_{\{\Phi^{\text{ext}}\}}$ with the lower index specifying the extended theory is more appropriate in this situation. Because the modular $S$ transformation exchanges the role of space and time, one can obtain the following mutual mapping,
@@
Hence by applying the property $S^{2}\sim 1$ of modular transformation, one can obtain the following expression,
@@
To distinguish its applications to the chiral characters, we denote the modular transformation as $\mathcal{S}$. Whereas the expression $\mathcal{D}$ produces the extended theory, the application $\mathcal{Q}$ does not provide any notrivial objects without introducing $\mathcal{S}_{\{\Phi\}^{\text{ext}}}$. The objects on the righthand side are completely determined by the algebraic relations. Hence one can apply an arbitrary number of defects and symmetries.
@@ Appendix B Realization of group extension corresponding to anomaly matching in lattice model
Figure 11: The construction of nonchiral gauging is implemented by gauging, inversion, and splitting.
@@
  • Applying the chiral Fradkin-Kadanoff transformation to a spin chain $(1+2)$ with the site labelled by $l=1,2,...,2L$. After the Fradkin-Kadanoff transformation, one obtains the site $l_{\text{PF}}=1,2,...,4L$.
@@
Because of the nonlocal property of the Fradkin-Kadanoff transformation[335], the coupled parafermion chains are strongly interacting after the transformations even when starting from a local Hamiltonian in the original paraspin representation. This property comes from the parafermi statistics of the chains. As the simplest example, we note the results for the $Z_{3}$ parafermion corresponding to the state Potts model.
$Z_{0}=\sum_{i^{(1)},i^{(2)}\in\{I,\epsilon\}^{2}}|\sum_{p=0}^{2}\chi^{(1)}_{i_{(1)},p}\overline{\chi}^{(2)}_{\overline{i}_{(2)},p}|^{2},$ (108)
$Z_{2/3}=\sum_{i^{(1)},i^{(2)}\in\{I,\epsilon\}^{2}}|\sum_{p=0}^{2}\chi^{(1)}_{i_{(1)},p+1}\overline{\chi}^{(2)}_{\overline{i}{(2)},p}|^{2},$ (109)
$Z_{1/3}=\sum_{i^{(1)},i^{(2)}\in\{I,\epsilon\}^{2}}|\sum_{p=0}^{2}\chi^{(1)}_{i_{(1)},p+2}\overline{\chi}^{(2)}_{\overline{i}{(2)},p}|^{2},$ (110)
@@
Construction of the BCFT for our models is an interesting problem, but one can obtain the corresponding OECFT by applying the method in [208] and the main text. By introducing the parity and charge for the simple current $J^{(1)}\overline{J}^{(2)}$ and studying the $Z_{N}$ extension of the spherical fusion category corresponding to the original bosonic CFT $(1)+(2)$ (or UV×$\times$×IR in the main text) and by taking a particular subalgebra called bulk semionization, one can obtain the algebraic data corresponding to the Cardy’s condition.
@@
The forms are the same as those by the author’s other work [32] and the $\{\Psi\}$ in the main text. Hence by identifying the respective object as $\{\phi^{(1)}\}\otimes\{\overline{\phi}^{(2)}\}$ or its complex conjugate, one can obtain the resultant fusion rule of OECFT.
@@
where $e^{-Ni\sqrt{q}\varphi(\infty)}$ is the so-called background charge or label of edge modes.The quasihole operator without matter is $e^{i\frac{r}{\sqrt{q}}\varphi}$ where $r=0,1,...q-1$. The above wavefunction is known as Laughlin’s wavefunction[349].
@@
One can apply the above analysis to general $Z_{N}$ models only by replacing some symbols in the existing literature (see review [143], and the general construction in [54]). This can include the case where the matter CFTs are $Z_{N}$ anomalous, but coupling this anomalous matter to flux, one can obtain the modified anomaly-free theory by modifying $J_{\text{tot}}$ to $J_{\text{tot}}\times(\text{flux})$. Hence the total partition function is still in the form Eq. (65) with the modified total charge zero. Moreover, contrary to the case with anomaly-free matter, the theory does not contain mock-modular forms. However, it has been pointed out that the resultant partition function is not in the Moore-Seiberg data[45]. Hence there exists difficulty to obtain bulk topological degeneracies from the existing framework in this case because of the close relationship between Moore-Seiberg data and bulk topological orders. Related problems have been summarized in the author’s other works [61, 73, 12]. Hence, there exists the following exchange relation
{Anomaly free matter= mock modular form}{Anomalous matter= Non $\begin{split}\{\text{Anomaly free matter= mock modular form}\}\\ \leftrightarrow\{\text{Anomalous matter= Non Moore-Seiberg}\}\end{split}$ (123)
@@ Appendix E Bulk and boundary RG flow: Application of Graham-Watts method
where $\alpha$ and $\gamma$ are the primary states labeling the UV and IR Cardy states and the matrix $D_{\text{RG}}^{\text{def}}$ is the matrix representation of the RG domain wall (We demonstrate the calculations elsewhere). The most fundamental point of this equation is the righthand side can become nonCardy states (such as Graham-Watts states[206, 208]) because the boundary $g$-value can show nondecreasing under bulk and boundary RG flow[305]. Similar application of the charged domain wall to boundary states is an interesting future problem.

### v3

@@ VI Acknowledgement
We thank Rui-Dong Zhu for introducing the work [104] to us about six years ago and Yuji Tachikawa and Yunqin Zheng for a fruitful collaboration closely related to new BCFTs in this manuscript. We also thank Urei Miura, Shumpei Iino, Kohki Kawabata, Tokiro Numasawa, Steven Simon, Masahiko G. Yamada, and Yuan Yao for the interesting discussions. We thank Bo Yang and Guangyue Ji for the related collaborations and for notifying us of the recent progress on the studies of the HLL in FQHEs and for teaching us the basic notions in FQHE through their lectures and tutorials. We also thank Hao Gang for notifying the work[119] for us. We acknowledge the support from NCTS and CTC-NTHU.
@@ Appendix A Insertion of topological defect and noninvertible chiral gauging
However, it is known that the $\mathcal{D}_{\psi}\sigma\overline{\sigma}$ corresponds to the disorder operator. This has been established in the study of fermionic models and this counting is known as fermionic zero modes. Hence, the interpretation of objects after the contraction needs further careful discussion, whereas they provide the correct partition functions.
In the above discussions, one can see the correspondence between $\phi$ and $\mathcal{D}$ evidently. We stress that the formal chiral and antichiral decomposition of the bulk fields $\Phi$ is not a tensor product. This is called the Deligne product [176] verified by “abstract nonsense" terminologically, but still, the formal decomposition is correct up to the chiral and antichiral decomposition of the partition function. Hence, one can trust the discussion on the amplitude in Section. II. As a summary, we note the relevant set of objects and identifications,
@@
The set $\{\mathcal{D}_{\alpha}\}$ corresponds to the chiral twist following modern literature on Fradkin-Kadanoff transformations[340]. Hence, one can implement the antichiral twist under the following identification,
@@
In successive discussions, we introduce the modular $S$ transformation to study the symmetry $\{\mathcal{Q}\}$. Here, we assume that the modular $S$ transformation does not change the total Hilbert space. Hence, the modular $S$ transformation is a mapping $S:\{\phi\}\times\{\overline{\phi}\}\rightarrow\{\Phi\}$. This can be achieved by constructing the modular $S$ matrix for the chiral and antichiral fields and applying the projection operator to the original Hilbert space. In other words, the construction of the original theory is fundamentally important. This choice of projection may become nontrivial when assuming the existence of extended chiral objects. For example, if one studies an orbifolded model, such as the $D$ series model, the object in the $A$ series model appears as such. Moreover, in the recent studies on Majorana and the parafermionic model, this extension of the Verlinde line also appears. Hence, the representation $S_{\{\Phi^{\text{ext}}\}}$ with the lower index specifying the extended theory is more appropriate in this situation. Because the modular $S$ transformation exchanges the role of space and time, one can obtain the following mutual mapping,
@@
Hence, by applying the property $S^{2}\sim 1$ of modular transformation, one can obtain the following expression,
@@
To distinguish its applications to the chiral characters, we denote the modular transformation as $\mathcal{S}$. Whereas the expression $\mathcal{D}$ produces the extended theory, the application $\mathcal{Q}$ does not provide any notrivial objects without introducing $\mathcal{S}_{\{\Phi\}^{\text{ext}}}$. The objects on the righthand side are completely determined by the algebraic relations. Hence, one can apply an arbitrary number of defects and symmetries.
@@ Appendix B Realization of group extension corresponding to anomaly matching in lattice model
Figure 11: The construction of nonchiral gauging is implemented by gauging, inversion, and splitting. By applying Fradkin-Kadanoff transformations to coupled spin chains labelled by $(1)$ and $(2)$ and space inversion operation only to spin chain (2), one can obtain the new composite $Z_{N}$ particle models. It should be noted that the space inversion is not included in a conformal transformations.
@@
  • Applying the chiral Fradkin-Kadanoff transformation to a spin chain $(1+2)$ with the site labelled by l=1,2,,2Ll=1,2,...,2L. After the Fradkin-Kadanoff transformation, one obtains the site $l_{\text{PF}}=1,2,...,4L$.
@@
Because of the nonlocal property of the Fradkin-Kadanoff transformation[340], the coupled parafermion chains are strongly interacting after the transformations even when starting from a local Hamiltonian in the original paraspin representation. This property comes from the parafermi statistics of the chains. As the simplest example, we note the results for the $Z_{3}$ parafermion corresponding to the coupled three state Potts model.
$Z_{0}=\sum_{i^{(1)},i^{(2)}\in\{I,\epsilon\}^{2}}|\sum_{p=0}^{2}\chi^{(1)}_{i^{(1)},p}\overline{\chi}^{(2)}_{\overline{i}^{(2)},p}|^{2},$ (108)
$Z_{2/3}=\sum_{i^{(1)},i^{(2)}\in\{I,\epsilon\}^{2}}|\sum_{p=0}^{2}\chi^{(1)}_{i^{(1)},p+1}\overline{\chi}^{(2)}_{\overline{i}^{(2)},p}|^{2},$ (109)
$Z_{1/3}=\sum_{i^{(1)},i^{(2)}\in\{I,\epsilon\}^{2}}|\sum_{p=0}^{2}\chi^{(1)}_{i^{(1)},p+2}\overline{\chi}^{(2)}_{\overline{i}^{(2)},p}|^{2},$ (110)
@@
[…]
@@
The forms are the same as those by the author’s other work [32] and the $\{\Psi\}$ in the main text. Hence, by identifying the respective object as $\{\phi^{(1)}\}\otimes\{\overline{\phi}^{(2)}\}$ or its complex conjugate, one can obtain the resultant fusion rule of OECFT.
@@
where $e^{-Ni\sqrt{q}\varphi(\infty)}$ is the so-called background charge or label of edge modes.The quasihole operator without matter is $e^{i\frac{r}{\sqrt{q}}\varphi}$ where r=0,1,q1r=0,1,...q-1. The above wavefunction is known as Laughlin’s wavefunction[354].
@@
One can apply the above analysis to general $Z_{N}$ models only by replacing some symbols in the existing literature (see review [144], and the general construction in [54]). This can include the case where the matter CFTs are $Z_{N}$ anomalous, but coupling this anomalous matter to flux, one can obtain the modified anomaly-free theory by modifying $J_{\text{tot}}$ to $J_{\text{tot}}\times(\text{flux})$. Hence, the total partition function is still in the form Eq. (65) with the modified total charge zero. Moreover, contrary to the case with anomaly-free matter, the theory does not contain mock-modular forms. However, it has been pointed out that the resultant partition function is not in the Moore-Seiberg data[45]. Hence, there exists difficulty to obtain bulk topological degeneracies from the existing framework in this case because of the close relationship between Moore-Seiberg data and bulk topological orders. Related problems have been summarized in the author’s other works [61, 73, 12]. Hence, there exists the following exchange relation
[…]
where $\Rightarrow$ represents coupling to the flux.
@@ Appendix E Bulk and boundary RG flow: Application of Graham-Watts method
where $\alpha$ and $\gamma$ are the primary states labeling the UV and IR Cardy states and the matrix $D_{\text{RG}}^{\text{def}}$ is the matrix representation of the RG domain wall (We demonstrate the calculations elsewhere). The most fundamental point of this equation is the righthand side can become nonCardy states (such as Graham-Watts states[210, 212]) because the boundary $g$-value can show nondecreasing under bulk and boundary RG flow[310]. Similar application of the charged domain wall to boundary states is an interesting future problem. We also note that the original motivation introducing RG domain wall has a close connection to the bulk and boundary RG flow[103].